two height are 54 and 30 on same base and arc on top of both. so find arc radius
1
Expert's answer
2013-06-05T07:49:28-0400
Theradius of an arc or segment is the radius
of the circle of which it is a part.
A formula is provided below for the radius given the width and height of the
arc.
Given an arc or segment with known width and height:
Radius = H/2 + W^2/8H
where:
W is the length of the chord defining the base of the arc
H is the height measured at the midpoint of the arc's base
So W = 54, H = 30
Radius = 30/2 + 54^2/8*30 = 27.15
Answer: 27.15